1. D(x2+2x-3)                  ans, 2x+2
  2. D2(xe3x-e4x)                 ans, 9xe3x+633x-16e4x
1
Expert's answer
2021-12-06T16:22:58-0500

"1. \ D(x^2+2x-3)=Dx^2+D(2x-3)=2x+2"


"2. \ D(xe^{3x}-e^{4x})=D(xe^{3x})-De^{4x}=Dx\cdot e^{3x}+x\cdot De^{3x}-De^{4x}=
\\
=e^{3x}+x\cdot e^{3x}\cdot D(3x)-e^{4x}\cdot D(4x)=e^{3x}+3xe^{3x}-4e^{4x}"


"D^2(xe^{3x}-e^{4x})=D((1+3x)e^{3x}-4e^{4x})=D(1+3x)\cdot e^{3x}+(1+3x)\cdot De^{3x}-4De^{4x}=3e^{3x}+(1+3x)\cdot 3e^{3x}-16e^{4x}=9xe^{3x}+6e^{3x}-16e^{4x}"


Answers: "1) 2x+2;\ \ \ 2) \ 9xe^{3x}+6e^{3x}-16e^{4x}."